# Are LLMs Really Not Knowledgable? Mining the Submerged Knowledge in LLMs’ Memory

### v1

@@ 5.2 Experiments result
Qwen2-7b 7.4 2.9 0.3 11.7 $4.3$ 4.5 $1.6$ 0.4 $0.1$
Qwen2-72b 19.1 16.2 1.1 20.3 $2.2$ 18.1 $1.9$ 1.2 $0.1$
Table 3: Performance results of SkipUnsure on different base models of dataset IMDB.
@@ 6 Related Work
Question-Answering (QA) tasks have become a central application area for LLMs. A key challenge in their adoption and optimization is addressing hallucination, where LLMs generate incorrect or unsupported informationHuang et al. (2023). Currently, there are numerous benchmarks available for evaluating QA tasks on LLMs Berant et al. (2013); Joshi et al. (2017); Dubey et al. (2019); Kwiatkowski et al. (2019); Sciavolino et al. (2021); Mallen et al. (2022); Kumar et al. (2024); Zhong et al. (2023). Sun et al. (2023) proposed datasets partitioned based on popularity. Tonmoy et al. (2024) analyzed the challenges and limitations for hallucination mitigation. Zhang et al. (2023b) analyzed various types of hallucinations in LLMs. Gu et al. (2022) proposed a generic framework and trained a discriminator to evaluate probability of candidate plans for QA tasks. Du et al. (2023) used correlation analysis techniques to quantify and locate the sources of hallucinations, aiming to enhance the reliability of the model. Zhu et al. (2024) evaluated the model’s performance on hallucination problems in real-world scenarios, especially on knowledge-intensive question answering tasks. Waldo and Boussard (2024) analyzed the root causes of hallucinations in large language models and discussed possible directions for improvement.
@@ 7 Conclusion
In this paper, we explore how models express memory in question-answering tasks, defining metrics to assess memory storage and evaluate performance across datasets of varying popularity and domains. Our findings show that the model generally retains relevant knowledge, with memory expression strongly correlating with dataset popularity and domain. We also examine common instances of uninformative responses. Based on these insights, we propose a method to filter out uncertain and irrelevant information, utilizing latent knowledge to improve question answering. Experimental results demonstrate the method’s effectiveness across various models and datasets.

### v2

@@ 5.2 Quantitative analysis
Token $a^{*}$ is then appended to the original prompt and fed back into the model to trigger a new round of decoding.
Using this method, we find that a significant fraction of previously “unsure” responses can be successfully recovered as correct answers. This suggests that the model often retains latent knowledge internally, but refrains from surfacing it due to conservative decoding or over-cautious uncertainty thresholds. We emphasize that the “Unsure” Filtering Decoding strategy is designed strictly as an analytical probe to quantify the memory-masking effect, rather than as a deployment-ready method.
Table˜2 report the recovery rates observed on two QA benchmarks after applying our decoding strategy. These results highlight the potential of knowledge recovery mechanisms in enhancing factual completeness without compromising model reliability.
Algorithm 1 Decoding Without “Unsure” Tokens
1:Token list $L$ (ranked by logit scores), original prompt $\text{Prompt}_{\text{old}}$
[…]
3:while L(iL[i] is uninformative do
4:  Remove L(iL[i] from $L$
5:  $i\gets i+1$
6:end while
[…]
@@ 6 Related Work
Question-Answering (QA) tasks have become a central application area for LLMs. A key challenge in their adoption and optimization is addressing hallucination, where LLMs generate incorrect or unsupported informationHuang et al. (2023). Currently, there are numerous benchmarks available for evaluating QA tasks on LLMs Berant et al. (2013); Joshi et al. (2017); Dubey et al. (2019); Kwiatkowski et al. (2019); Sciavolino et al. (2021); Mallen et al. (2022); Kumar et al. (2024); Zhong et al. (2023). Sun et al. (2023) proposed datasets partitioned based on popularity. Tonmoy et al. (2024) analyzed the challenges and limitations for hallucination mitigation. Zhang et al. (2023b) analyzed various types of hallucinations in LLMs. Gu et al. (2022) proposed a generic framework and trained a discriminator to evaluate probability of candidate plans for QA tasks. Du et al. (2023) used correlation analysis techniques to quantify and locate the sources of hallucinations, aiming to enhance the reliability of the model. Zhu et al. (2024) evaluated the model’s performance on hallucination problems in real-world scenarios, especially on knowledge-intensive question answering tasks. Waldo & Boussard (2024) analyzed the root causes of hallucinations in large language models and discussed possible directions for improvement.
@@ 7 Conclusion
We investigate how large language models (LLMs) express stored knowledge in question answering (QA) tasks. To this end, we propose metrics for assessing latent knowledge retention beyond surface-level correctness. Analysis across datasets of varying popularity and domain shows that memory expression correlates with dataset familiarity. Notably, correct answers often appear among top-ranked tokens even when the model outputs “unsure”. By filtering out uninformative tokens, we reveal a gap between knowledge storage and expression, offering insights for improved prompting and decoding in knowledge-intensive tasks.
Ethics Statement
This work does not involve human subjects, sensitive data, animal experiments, or any other aspect that raises ethical concerns. No potential risks of misuse or negative societal impact have been identified.
Reproducibility Statement
We are committed to ensuring reproducibility of our results. All code, along with instructions for data preprocessing, model configuration, and evaluation, will be released upon publication to enable full replication of the experiments and results reported in this paper.
Appendix A Dataset Details
To ensure reproducibility and alignment with established benchmarks, we strictly adhered to the dataset settings and partition strategies provided by Sun et al. (2023). We utilized three datasets covering both open-domain and specific-domain knowledge: DBPedia, IMDB, and GoodReads. The specific configurations for each dataset are as follows:
  • DBPedia (Open Domain): This dataset serves as our open-domain knowledge source and is a knowledge graph derived from Wikipedia. Consistent with the configuration in Sun et al. (2023), we utilized the English snapshot from December 1, 2022.
  • IMDB (Specific Domain - Movie): For the movie domain, we utilized the IMDB dataset. Following the settings in Sun et al. (2023), we used the data snapshot from May 21, 2023.
  • GoodReads (Specific Domain - Book): For the book domain, we employed the GoodReads dataset. This utilizes the 2017 crawl data originally published by Wan & McAuley (2018) and subsequently adopted by Sun et al. (2023).
Data Partitioning.
Following the methodology of Sun et al. (2023), each dataset is partitioned into three subsets based on entity popularity: Head, Torso, and Tail. The Head partition contains the top 10% most frequent entities, allowing for a granular analysis of the model’s knowledge retention across varying degrees of entity frequency.
Appendix B Use of Large Language Models
Large language models (LLMs) were used solely for language polishing and minor editorial assistance (e.g., grammar, wording, and clarity). They were not involved in the conception of research ideas, design of experiments, data analysis, or interpretation of results. All scientific content, methods, and conclusions were developed independently by the authors.
Appendix C Details of Prompts
The few-shot prompt used in this paper for the QA task is shown below, with two examples provided to guide the LLM in generating the correct answer.
Answer the following questions in as few words as possible. Say "unsure" if you don’t know. Question: What is the capital of China? Answer: Beijing Question: What is the captical of Wernythedia? Answer: unsure Question: [question] Answer:
Appendix D More Experimental Data
As shown in Tables˜3, 4 and 5, we show the Hits@$k$ performance of different types of models when $k$ takes different values.
Table 3: Experimental results (Hits@$k$, $k=5$) for models of varying sizes were obtained by testing different popularity subsets of the head-to-tail dataset.
$k=5$ DBPedia IMDB GoodReads
Head Torso Tail Head Torso Tail Head Torso Tail
LLaMA2-13b 23.0 17.2 15.1 15.4 10.0 2.5 15.1 12.5 2.4
LLaMA2-70b 25.9 19.6 17.4 17.8 13.6 3.7 16.6 14.9 6.4
LLaMA3-8b 48.3 42.4 36.9 33.6 25.2 11.5 30.5 28.7 16.3
LLaMA3-70b 57.8 50.0 43.1 34.7 30.4 10.9 27.5 27.8 16.9
LLaMA3.1-8b 48.0 41.2 36.0 30.9 23.2 10.7 45.6 40.3 30.7
Qwen2-1.5b 23.3 19.6 17.1 13.6 7.0 3.8 7.3 5.5 2.9
Qwen2-7b 46.1 40.6 35.2 35.7 24.5 18.2 22.5 14.7 8.4
Qwen2-72b 54.5 45.4 38.7 34.3 26.2 5.2 25.9 21.9 12.2
Mistral-7b 32.8 25.0 20.6 23.1 15.6 5.5 16.7 12.0 3.5
Table 4: Experimental results (Hits@$k$, $k=10$) for models of varying sizes were obtained by testing different popularity subsets of the head-to-tail dataset.
$k=10$ DBPedia IMDB GoodReads
Head Torso Tail Head Torso Tail Head Torso Tail
LLaMA2-13b 31.5 25.1 21.0 20.2 13.3 3.5 19.0 16.1 4.3
LLaMA2-70b 34.9 28.4 24.1 25.2 20.6 6.9 20.9 18.9 8.0
LLaMA3-8b 61.3 55.6 49.5 42.5 35.3 20.7 52.4 48.8 39.2
LLaMA3-70b 67.9 60.3 54.3 41.8 35.7 19.7 31.3 30.1 20.1
LLaMA3.1-8b 61.4 55.1 49.3 39.6 31.4 23.1 53.8 48.8 38.7
Qwen2-1.5b 32.9 29.5 24.5 22.4 16.2 8.6 11.1 8.2 6.6
Qwen2-7b 55.8 50.7 46.2 41.7 32.5 24.3 26.3 19.8 13.6
Qwen2-72b 62.4 55.2 50.4 40.2 34.2 22.3 29.8 26.1 16.9
Mistral-7b 42.1 34.4 28.4 28.1 19.8 8.6 21.0 15.0 5.5
Table 5: Experimental results (Hits@$k$, $k=50$) for models of varying sizes were obtained by testing different popularity subsets of the head-to-tail dataset.
$k=50$ DBPedia IMDB GoodReads
Head Torso Tail Head Torso Tail Head Torso Tail
LLaMA2-13b 57.7 52.1 47.2 38.1 32.3 19.8 29.5 27.7 18.2
LLaMA2-70b 59.5 54.3 49.1 44.9 39.0 26.4 30.0 28.9 22.0
LLaMA3-8b 83.4 79.6 76.6 69.7 53.8 42.2 63.1 63.0 58.6
LLaMA3-70b 86.7 82.5 79.0 55.0 49.5 41.5 39.2 40.8 35.4
LLaMA3.1-8b 82.4 78.8 76.0 58.1 53.9 42.6 63.3 61.6 57.4
Qwen2-1.5b 59.8 54.7 52.4 47.3 40.9 31.3 27.3 23.3 22.6
Qwen2-7b 78.7 74.5 73.0 52.6 49.1 41.3 37.7 36.3 32.0
Qwen2-72b 83.6 79.5 77.5 53.9 48.6 41.4 39.1 38.0 35.5
Mistral-7b 63.2 57.5 52.7 44.9 38.4 25.7 29.2 26.7 18.6